OpenStudy (firejay5):

Find the midpoint of each side of the right triangle with vertices: D: (-1, 2) E: (8, 2) F: (8, 14) Medal will be rewarded for your patronage! :D

OpenStudy (firejay5):

@satellite73 @zepdrix @ganeshie8

OpenStudy (anonymous):

midpoint of E and F is fairly clearly \((8,8)\) it is the average of each coordinate add up and divide by 2

OpenStudy (anonymous):

midpoint of \( (-1, 2), (8, 2)\) is \((\frac{8-1}{2},\frac{2+2}{2})=(\frac{7}{2},2)\)

OpenStudy (firejay5):

so how is E & F (8,8)

OpenStudy (anonymous):

the average of 2 and 14 is 8, i.e. \(\frac{2+14}{2}=\frac{16}{2}=8\)

OpenStudy (anonymous):

as for the first coordinate, both numbers are the same, so that is a vertical line or you can think that the average of 8 and 8 is 8

OpenStudy (firejay5):

why does 8 + -1 = 9 / 2 = 9/2 why is it 8 - 1/2

OpenStudy (anonymous):

i am pretty sure i didn't write either of those

OpenStudy (firejay5):

\[\frac{ (8 - 1) }{ 2 } + \frac{ (2 + 2) }{ 2 } = (\frac{ 7 }{ 2 }, 2)\]

OpenStudy (firejay5):

look at the third comment at the beginning

OpenStudy (anonymous):

the average of 8 and -1 is \(\frac{7}{2}\)

OpenStudy (anonymous):

and the average of 2 and 2 is 2

OpenStudy (firejay5):

you just forgot to add the "+" an but -1 ---> (-1)